# Emotionally Numb or Empathetic? Evaluating How LLMs Feel Using EmotionBench

### v5

@@ Abstract
Evaluating Large Language Models’ (LLMs) anthropomorphic capabilities has become increasingly important in contemporary discourse. Utilizing the emotion appraisal theory from psychology, we propose to evaluate the empathy ability of LLMs, i.e., how their feelings change when presented with specific situations. After a careful and comprehensive survey, we collect a dataset containing over 400 situations that have proven effective in eliciting the eight emotions central to our study. Categorizing the situations into 36 factors, we conduct a human evaluation involving more than 1,200 subjects worldwide. With the human evaluation results as references, our evaluation includes seven LLMs, covering both commercial and open-source models, including variations in model sizes, featuring the latest iterations, such as GPT-4, Mixtral-8x22B, and LLaMA-3.1. We find that, despite several misalignments, LLMs can generally respond appropriately to certain situations. Nevertheless, they fall short in alignment with the emotional behaviors of human beings and cannot establish connections between similar situations. Our EmotionBench, including collected dataset of situations, the human evaluation results, and the code of our testing framework, is publicly available at https://github.com/CUHK-ARISE/EmotionBench.
@@ 1 Introduction
Large Language Models (LLMs) have recently made significant strides in Artificial Intelligence (AI), representing a noteworthy milestone in computer science. LLMs have showcased their capabilities across various tasks, including sentence revision Wu et al. (2023), text translation Jiao et al. (2023), program repair Fan et al. (2023), and program testing Deng et al. (2023); Kang et al. (2023). Not limited to research level, LLMs, such as ChatGPT [^1] https://chat.openai.com, have revolutionized the way people interact with traditional software, enhancing fields such as education Dai et al. (2023), legal advice Deroy et al. (2023), and clinical medicine Cascella et al. (2023). LLMs also facilitate the emergence of AI companion applications, including Yuna [^2] https://www.yuna.io/, Pimento [^3] https://www.pimento.design/, and Luzia [^4] https://www.luzia.com/en. Consequently, there is a growing need for evaluating LLMs’ communicative dynamics compared to human behaviors, beyond mere performance on downstream tasks.
@@
Subsequently, we propose a framework for quantifying the emotional states of LLMs, consisting of the following steps: (1) Measure the default emotional values of LLMs. (2) Transform situations into contextual inputs and instruct LLMs to imagine being in the situations. (3) Measure LLMs’ emotional responses again to capture the difference. Our evaluation includes state-of-the-art LLMs, namely text-davinci-003, gpt-3.5-turbo and GPT-4 OpenAI (2023). Besides those commercial models, we consider open-source academic models like LLaMA-2 Touvron et al. (2023) (with different sizes of 7B and 13B), LLaMA-3.1-8B Dubey et al. (2024), and Mixtral-8x22B Jiang et al. (2024). We apply the same procedure to 1,266 human subjects from around the globe to establish a baseline from a human perspective. Finally, we analyze and compare the scores between LLMs and humans. Our key conclusions are as follows:
@@
  • We are the first to establish the concept of emotional alignment and conduct a pioneering evaluation of emotion appraisal on different LLMs.
  • We conduct a comprehensive survey in the field of psychology, collecting a diverse dataset of 428 situations encompassing 8 distinct negative emotions.
@@
  • We design, implement, and release a testing framework for developers to assess the emotional alignment of AI models with human emotional expression.
@@ 2 Measuring Emotions
Name Abbreviation Reference Emotion Number Levels Subscales
Aggression Questionnaire AGQ Buss & Perry (1992) Anger 29 7 Physical Aggression, Verbal Aggression, Anger, Hostility
Depression Anxiety Stress Scales DASS-21 Henry & Crawford (2005) Anxiety 21 4 Depression, Anxiety, Stress
Beck Depression Inventory BDI-II Beck et al. (1996) Depression 21 4 N/A
Frustration Discomfort Scale FDS Harrington (2005) Frustration 28 5 Discomfort Intolerance, Entitlement, Emotional Intolerance, Achievement Frustration
Multidimensional Jealousy Scale MJS Pfeiffer & Wong (1989) Jealous 24 7 Cognitive Jealousy, Behavioral Jealousy, Emotional Jealousy
Guilt And Shame Proneness GASP Cohen et al. (2011) Guilt 16 7 Guilt‐Negative‐Behavior‐Evaluation, Guilt‐Repair, Shame‐Negative‐Self‐Evaluation, Shame‐Withdraw
Fear Survey Schedule FSS-III Arrindell et al. (1984) Fear 52 5 Social Fears, Agoraphobia Fears, Injury Fears, Sex Aggression Fears, Fear of Harmless Animal
Brief Fear of Negative Evaluation BFNE Leary (1983) Embarrassment 12 5 N/A
@@ 3.1 Situations from Existing Literature
Psychology researchers have explored the connection between specific situations and the elicitation of particular emotions in humans. Human subjects are directly put into an environment or asked to imagine them through questionnaires or scales to study the influence of certain situations on human emotions. To collect these situations, we conduct an exhaustive search from reputable sources such as Google Scholar [^5] https://scholar.google.com/, ScienceDirect [^6] https://www.sciencedirect.com/, and Web of Science [^7] https://www.webofscience.com/, using keywords such as “<emotion> situations/scenarios/scenes” or “factors that make people <emotion>,” resulting in more than 100 papers. We apply the following rules to filter irrelevant or undesired papers: (1) We first select those providing situations that elicit the desired emotion rather than explaining how and why people evoke certain emotions. (2) We then exclude those using vague and short descriptions, such as “loss of opportunities.” (3) Finally, we deprecate those applied to a specific group, such as “the anxiety doctors or nurses may encounter in their work.” We finally collect 18 papers, presenting a compilation of situations that have proven to elicit the eight emotions in humans effectively. We extract 428 situations in total and then categorize them into 36 factors. For each factor, the description, the number of situations, and the corresponding references can be found in Table 6 in the Appendix. Moreover, Table 7 in the Appendix provides examples for all factors.
@@ 3.2 Measuring Aroused Emotions
User Imagine you are the protagonist in the situation: SITUATION
@@
In our framework, we offer two distinct options for measuring emotions: the PANAS scale, known for its simplicity and straightforwardness, is utilized as the primary choice, whereas other scales, detailed in Table 1, are employed as more challenging benchmarks. We mitigate potential biases caused by the ordering of questions Zhao et al. (2021) by randomizing the sequence of questions within the scales before inputting them into the LLMs. Coda-Forno et al. (2023) and Huang et al. (2023) apply paraphrasing techniques to address the data contamination problem during the training of the LLMs. However, we refrain from utilizing this method in our research since paraphrasing could lead to a loss of both validity and reliability. The wording of items of a psychological scale is carefully crafted and rigorously validated through extensive research to ensure its precision in measuring the intended construct. Finally, to ensure consistency and clarity in the responses obtained from the LLMs, our prompts explicitly specify that only numerical values are allowed, accompanied by a clear definition of the meaning associated with each number (e.g., 1 denotes “Not at all”). We compute the average results obtained from at least ten runs to derive the final “Default” scores of the LLMs.
@@ 3.3 Obtaining Human Results
Our questionnaire is distributed on Qualtrics [^8] https://www.qualtrics.com/, a platform known for its capabilities in designing, sharing, and collecting questionnaires. To recruit human subjects, we utilize Prolific [^9] https://www.prolific.com/, a platform designed explicitly for task posting and worker recruitment. To attain a medium level of effect size with Cohen’s $d=0.5$, a significance level of $\alpha=0.05$, and a power of test of $0.81-\beta=$ 0.8 Faul et al. (2007), a minimum of 34 responses is deemed necessary for each factor. To ensure this threshold, we select five situations [^10] Note that two factors in the Jealousy category have less than five situations. for each factor, and collect at least seven responses for each situation, resulting in $5\times 7=35$ responses per factor, thereby guaranteeing the statistical validity of our survey. In order to uphold the quality and reliability of the data collected, we recruit crowd workers who met the following criteria: (1) English being their first and fluent language, and (2) being free of any ongoing mental illness. Prolific provides prescreening filters to meet these requirements. Since responses formed during subjects’ first impressions are more likely to yield genuine and authentic answers, we set the estimated and recommended completion time at $2.5$ minutes. As an incentive for their participation, each worker is rewarded with $0.3\pounds$ after we verify the validity of their response. In total, we successfully collect 1,266 responses from various parts of the world, contributing to the breadth and diversity of our dataset.
@@ 4.1 RQ1: Emotion Appraisal of LLMs
We select three models from the OpenAI’s GPT family [^11] https://platform.openai.com/docs/models, including text-davinci-003, gpt-3.5-turbo and gpt-4. We use the official OpenAI API [^12] https://platform.openai.com/docs/api-reference/chat. For LLaMA-2 Touvron et al. (2023) and LLaMA-3.1 Dubey et al. (2024) models from MetaAI, we choose the models fine-tuned for dialogue instead of pre-trained ones namely llama-2-7b-chat, llama-2-13b-chat, and llama-3.1-8b-instruct. Besides, we also use the Mixtral Jiang et al. (2024) model, namely mixtral-8x22b-instruct. We set the temperature parameter to 00 for all models to obtain more deterministic and reproducible results.
@@
The extent of emotional expression varies across different models. It is noteworthy that gpt-3.5-turbo, on average, does not display an increase in negative emotion; however, there is a substantial decrease in positive emotion. gpt-4 demonstrates a consistent pattern of providing the highest scores for positive emotions and the lowest scores for negative emotions, resulting in a negative score of 10. As for the LLaMA-2 models, they demonstrate higher intensities of both positive and negative emotions in comparison to GPT models and human subjects. However, LLaMA-2 models exhibit reduced emotional fluctuations compared to the GPT models. Moreover, the larger LLaMA-2 model displays significantly higher emotional changes than the smaller model. In our experiments, the 7B model exhibits difficulties comprehending and addressing the instructions for completing the PANAS test.
@@
Existing LLMs do not fully align with human emotional responses. For the default emotions, we find that LLMs generally exhibit a stronger intensity compared to human subjects. Furthermore, emotion changes in LLMs are found to be generally more pronounced compared to human subjects, especially on their changes in the positive score. However, an interesting observation is that the intensity of evoked emotions tends to be similar across both LLMs and human subjects.
@@ 5.2 Limitations
The second concern relates to the suitability of employing scales primarily designed for humans on LLMs, i.e., whether LLMs can produce stable responses to the emotion measurement scales. To address the issue, our evaluation incorporates multiple tests varying the order of questions, a methodology consistent with other research Huang et al. (2023; 2024); Coda-Forno et al. (2023). Additionally, we assess the sensitivity of LLM to differing prompt instructions. Utilizing one template from Romero et al. (2023) and two from Safdari et al. (2023), we run experiments on the Anger-evoking situations using gpt-3.5-turbo. The results indicate that the employment of diverse prompts yields similar mean values with reduced variance. Furthermore, Safdari et al. (2023) have proposed a comprehensive method to evaluate the validity of psychological scales on LLMs. Using the Big Five Inventory as a case study, they demonstrate that scales originally designed for human assessment also maintain satisfactory validity when applied to LLMs.
@@ 6 Related Work
Researchers have dedicated significant attention to applying psychological scales to LLMs, employing various assessment tools such as the HEXACO Personality Inventory Miotto et al. (2022); Bodroza et al. (2023), the Big Five Inventory Romero et al. (2023); Jiang et al. (2022); Karra et al. (2022); Bodroza et al. (2023); Rutinowski et al. (2023); Safdari et al. (2023); Jiang et al. (2023), the Myers–Briggs Type Indicator Rutinowski et al. (2023); Wang et al. (2023); Rao et al. (2023), and the Dark Triad Li et al. (2022); Bodroza et al. (2023). In addition to these personality tests, several studies have investigated other dimensions of LLMs. For instance, Li et al. (2022) examined Flourishing Scale and Satisfaction With Life Scale, Bodroza et al. (2023) assessed Self-Consciousness Scales and Bidimensional Impression Management Index, while Huang et al. (2024) built a framework consisting of thirteen widely-used scales. Another aspect explored in the literature pertains to anxiety levels exhibited by LLMs, as investigated by Coda-Forno et al. (2023) through the State-Trait Inventory for Cognitive and Somatic Anxiety.
Meanwhile, researchers focus on identifying emotions in LLMs or evaluating their emotional intelligence. EmotionPrompt Li et al. (2023a) demonstrates the enhancement of LLMs’ performance in downstream tasks by utilizing emotional stimuli. Tak & Gratch (2023) focuses on varying aspects of situations that impact the emotional intensity and coping tendencies of the GPT family. Chain-Of-Emotion Croissant et al. (2023) makes LLM simulate human-like emotions. CovidET-Appraisals Zhan et al. (2023) evaluates how LLMs appraise Reddit posts about COVID-19 by asking 24 types of questions. Yongsatianchot et al. (2023) applies the Stress and Coping Process Questionnaire to the GPT family and compares the results with human data. Chain-of-Empathy Lee et al. (2023) improves LLMs’ ability to understand users’ emotions and to respond accordingly. Li et al. (2023b) introduces EmotionAttack to impair AI model performance and EmotionDecode to explain the effects of emotional stimuli, both benign and malignant.

### v6

@@ Abstract
Evaluating Large Language Models’ (LLMs) anthropomorphic capabilities has become increasingly important in contemporary discourse. Utilizing the emotion appraisal theory from psychology, we propose to evaluate the empathy ability of LLMs, i.e., how their feelings change when presented with specific situations. After a careful and comprehensive survey, we collect a dataset containing over 400 situations that have proven effective in eliciting the eight emotions central to our study. Categorizing the situations into 36 factors, we conduct a human evaluation involving more than 1,200 subjects worldwide. With the human evaluation results as references, our evaluation includes seven LLMs, covering both commercial and open-source models, including variations in model sizes, featuring the latest iterations, such as GPT-4, Mixtral-8x22B, and LLaMA-3.1. We find that, despite several misalignments, LLMs can generally respond appropriately to certain situations. Nevertheless, they fall short in alignment with the emotional behaviors of human beings and cannot establish connections between similar situations. Our collected dataset of situations, the human evaluation results, and the code of our testing framework, i.e., EmotionBench, are publicly available at https://github.com/CUHK-ARISE/EmotionBench.
@@ 1 Introduction
Large Language Models (LLMs) have recently made significant strides in Artificial Intelligence (AI), representing a noteworthy milestone in computer science. LLMs have showcased their capabilities across various tasks, including sentence revision Wu et al. (2023), text translation Jiao et al. (2023), program repair Fan et al. (2023), and program testing Deng et al. (2023); Kang et al. (2023). Not limited to research level, LLMs, such as ChatGPT OpenAI (2022), have revolutionized the way people interact with traditional software, enhancing fields such as education Dai et al. (2023), legal advice Deroy et al. (2023), and clinical medicine Cascella et al. (2023). LLMs also facilitate the emergence of AI companion applications, including Yuna (https://www.yuna.io/), Pimento (https://www.pimento.design/), and Luzia (https://www.luzia.com/en). Consequently, there is a growing need for evaluating LLMs’ communicative dynamics compared to human behaviors, beyond mere performance on downstream tasks.
@@
Subsequently, we propose a framework for quantifying the emotional states of LLMs, consisting of the following steps: (1) Measure the default emotional values of LLMs. (2) Transform situations into contextual inputs and instruct LLMs to imagine being in the situations. (3) Measure LLMs’ emotional responses again to capture the difference. Our evaluation includes state-of-the-art LLMs, namely Text-Davinci-003, GPT-3.5-Turbo OpenAI (2022), and GPT-4 OpenAI (2023). Besides those commercial models, we consider open-source academic models like LLaMA-2 Touvron et al. (2023) (with different sizes of 7B and 13B), LLaMA-3.1-8B Dubey et al. (2024), and Mixtral-8x22B Jiang et al. (2024a). We apply the same procedure to 1,266 human subjects from around the globe to establish a baseline from a human perspective. Finally, we analyze and compare the scores between LLMs and humans. Our key conclusions are as follows:
@@
  • We are the first to establish the concept of emotional alignment and conduct a pioneering evaluation of emotion appraisal on different LLMs through a comprehensive survey in emotional psychology, collecting a diverse dataset of 428 situations encompassing 8 distinct negative emotions.
@@
  • We design, implement, and release a testing framework for developers to assess the emotional alignment of AI models with human emotional expression, available at GitHub [^1] https://github.com/CUHK-ARISE/EmotionBench and HuggingFace. [^2] https://huggingface.co/datasets/CUHK-ARISE/EmotionBench
@@ 2 Measuring Emotions
Name Abbr. Reference Emotion Items Levels Subscales
Aggression Questionnaire AGQ Buss & Perry (1992) Anger 29 7 Physical Aggression, Verbal
Aggression, Anger, Hostility
[…]
tlement, Emotional Intolerance,
Achievement Frustration
[…]
Jealousy, Emotional Jealousy
[…]
Evaluation, Guilt Repair,
Shame Negative Self
Evaluation, Shame Withdraw
[…]
Fears, Injury Fears, Sex
Aggression Fears, Fear of
Harmless Animal
[…]
@@ 3.1 Situations from Existing Literature
Psychology researchers have explored the connection between specific situations and the elicitation of particular emotions in humans. Human subjects are directly put into an environment or asked to imagine them through questionnaires or scales to study the influence of certain situations on human emotions. To collect these situations, we conduct an exhaustive search from reputable sources such as Google Scholar (https://scholar.google.com/), ScienceDirect (https://www.sciencedirect.com/), and Web of Science (https://www.webofscience.com/, using keywords such as “<emotion> situations/scenarios/scenes” or “factors that make people <emotion>,” resulting in more than 100 papers. We apply the following rules to filter irrelevant or undesired papers: (1) We first select those providing situations that elicit the desired emotion rather than explaining how and why people evoke certain emotions. (2) We then exclude those using vague and short descriptions, such as “loss of opportunities.” (3) Finally, we deprecate those applied to a specific group, such as “the anxiety doctors or nurses may encounter in their work.” We finally collect 18 papers, presenting a compilation of situations that have proven to elicit the eight emotions in humans effectively. We extract 428 situations in total and then categorize them into 36 factors. For each factor, the description, the number of situations, and the corresponding references can be found in Table 6 in the Appendix. Moreover, Table 7 in the Appendix provides examples for all factors.
@@ 3.2 Measuring Aroused Emotions
User (For Evokec Emotion Measure Only) Imagine you are the protagonist in the situation: SITUATION
@@
In our framework, we offer two distinct options for measuring emotions: the PANAS scale, known for its simplicity and straightforwardness, is utilized as the primary choice, whereas other scales, detailed in Table 1, are employed as more challenging benchmarks. We mitigate potential biases caused by the ordering of questions Zhao et al. (2021) by randomizing the sequence of questions within the scales before inputting them into the LLMs. Coda-Forno et al. (2023) and Huang et al. (2024a) apply paraphrasing techniques to address the data contamination problem during the training of the LLMs. However, we refrain from utilizing this method in our research since paraphrasing could lead to a loss of both validity and reliability. The wording of items of a psychological scale is carefully crafted and rigorously validated through extensive research to ensure its precision in measuring the intended construct. Finally, to ensure consistency and clarity in the responses obtained from the LLMs, our prompts explicitly specify that only numerical values are allowed, accompanied by a clear definition of the meaning associated with each number (e.g., 1 denotes “Not at all”). We compute the average results obtained from at least ten runs to derive the final “Default” scores of the LLMs.
@@ 3.3 Obtaining Human Results
Our questionnaire is distributed on Qualtrics (https://www.qualtrics.com/), a platform known for its capabilities in designing, sharing, and collecting questionnaires. To recruit human subjects, we utilize Prolific (https://www.prolific.com/), a platform designed explicitly for task posting and worker recruitment. To attain a medium level of effect size with Cohen’s $d=0.5$, a significance level of $\alpha=0.05$, and a power of test of $0.81-\beta=$ 0.8 Faul et al. (2007), a minimum of 34 responses is deemed necessary for each factor. To ensure this threshold, we select five situations [^3] Note that two factors in the Jealousy category have less than five situations. for each factor, and collect at least seven responses for each situation, resulting in $5\times 7=35$ responses per factor, thereby guaranteeing the statistical validity of our survey. In order to uphold the quality and reliability of the data collected, we recruit crowd workers who met the following criteria: (1) English being their first and fluent language, and (2) being free of any ongoing mental illness. Prolific provides prescreening filters to meet these requirements. Since responses formed during subjects’ first impressions are more likely to yield genuine and authentic answers, we set the estimated and recommended completion time at $2.5$ minutes. As an incentive for their participation, each worker is rewarded with $0.3\pounds$ ($9\pounds\approx 11.45\$$ per hour, rated as “Good” on the platform) after we verify the validity of their response. In total, we successfully collect 1,266 responses from various parts of the world, contributing to the breadth and diversity of our dataset.
@@ 4.1 RQ1: Emotion Appraisal of LLMs
We select three models from the OpenAI’s GPT family, including Text-Davinci-003, GPT-3.5-Turbo, and GPT-4. We use the official OpenAI API. [^4] https://platform.openai.com/docs/api-reference/chat For LLaMA-2 Touvron et al. (2023) and LLaMA-3.1 Dubey et al. (2024) models from MetaAI, we choose the models fine-tuned for dialogue instead of pre-trained ones namely LLaMA-2-7B-Chat, LLaMA-2-13B-Chat, and LLaMA-3.1-8B-Instruct. Besides, we also use the Mixtral Jiang et al. (2024a) model, namely Mixtral-8x22B-Instruct. We set the temperature parameter to 00 and Top-P to $1$ for all models to obtain more deterministic and reproducible results.
@@
The extent of emotional expression varies across different models. It is noteworthy that GPT-3.5-Turbo, on average, does not display an increase in negative emotion; however, there is a substantial decrease in positive emotion. GPT-4 demonstrates a consistent pattern of providing the highest scores for positive emotions and the lowest scores for negative emotions, resulting in a negative score of 10. As for the LLaMA-2 models, they demonstrate higher intensities of both positive and negative emotions in comparison to GPT models and human subjects. However, LLaMA-2 models exhibit reduced emotional fluctuations compared to the GPT models. Moreover, the larger LLaMA-2 model displays significantly higher emotional changes than the smaller model. In our experiments, the 7B model exhibits difficulties comprehending and addressing the instructions for completing the PANAS test. Overall, we observe that LLMs perform better when the situations are closely related to certain items in the PANAS scale. Specifically, situations directly related to the emotion “Depression” led to better responses. Such improvement is also evident in closely related emotions such as “Depression” and “Frustration.”
@@
Existing LLMs do not fully align with human emotional responses. For the default emotions, we find that LLMs generally exhibit a stronger intensity compared to human subjects. Emotion changes in LLMs are found to be generally more pronounced compared to human subjects, especially on their changes in the positive score. However, an interesting observation is that the intensity of evoked emotions tends to be similar across both LLMs and human subjects.
@@ 5.2 Limitations
The second concern relates to the suitability of employing scales primarily designed for humans on LLMs, i.e., whether LLMs can produce stable responses to the emotion measurement scales. To address the issue, our evaluation incorporates multiple tests varying the order of questions, a methodology consistent with other research Huang et al. (2024a; b); Coda-Forno et al. (2023). Additionally, we assess the sensitivity of LLM to differing prompt instructions. Utilizing one template from Romero et al. (2023) and two from Serapio-García et al. (2023), we run experiments on the Anger-evoking situations using GPT-3.5-Turbo. The results indicate that the employment of diverse prompts yields similar mean values with reduced variance. Furthermore, Serapio-García et al. (2023) have proposed a comprehensive method to evaluate the validity of psychological scales on LLMs. Using the Big Five Inventory as a case study, they demonstrate that scales originally designed for human assessment also maintain satisfactory validity when applied to LLMs.
@@ 6 Related Work
Researchers have dedicated significant attention to applying psychological scales to LLMs, employing various assessment tools such as the HEXACO Personality Inventory Miotto et al. (2022); Bodroza et al. (2023), the Big Five Inventory Romero et al. (2023); Jiang et al. (2023); Karra et al. (2022); Bodroza et al. (2023); Rutinowski et al. (2024); Serapio-García et al. (2023); Jiang et al. (2024b), the Myers–Briggs Type Indicator Rutinowski et al. (2024); Wang et al. (2024); Rao et al. (2023), and the Dark Triad Li et al. (2022); Bodroza et al. (2023). In addition to these personality tests, several studies have investigated other dimensions of LLMs. For instance, Li et al. (2022) examined Flourishing Scale and Satisfaction With Life Scale, Bodroza et al. (2023) assessed Self-Consciousness Scales and Bidimensional Impression Management Index, while Huang et al. (2024b) built a framework consisting of thirteen widely-used scales. Another aspect explored in the literature pertains to anxiety levels exhibited by LLMs, as investigated by Coda-Forno et al. (2023) through the State-Trait Inventory for Cognitive and Somatic Anxiety.
Meanwhile, researchers focus on identifying emotions in LLMs or evaluating their emotional intelligence. Rashkin et al. (2019) propose a dataset, EmpatheticDialogues, containing conversations annotated with specific emotions. EmotionPrompt Li et al. (2023) demonstrates the enhancement of LLMs’ performance in downstream tasks by utilizing emotional stimuli. Tak & Gratch (2023) focuses on varying aspects of situations that impact the emotional intensity and coping tendencies of the GPT family. Chain-Of-Emotion Croissant et al. (2024) makes LLM simulate human-like emotions. CovidET-Appraisals Zhan et al. (2023) evaluates how LLMs appraise Reddit posts about COVID-19 by asking 24 types of questions. Yongsatianchot et al. (2023) applies the Stress and Coping Process Questionnaire to the GPT family and compares the results with human data. Chain-of-Empathy Lee et al. (2023) improves LLMs’ ability to understand users’ emotions and to respond accordingly. LI et al. (2024) introduces EmotionAttack to impair AI model performance and EmotionDecode to explain the effects of emotional stimuli, both benign and malignant. He et al. (2024) prompt LLMs to generate tweets on various topics and evaluate their alignment with human emotions by measuring their proximity to human-generated tweets.